Astronomy Report – January 15th

Science – Astronomy

Name: Cynthia Fuertes Panizo

Crew: 187

Date: 15JAN2018

Sky Conditions: N/A

Wind Conditions: N/A

Observation Start Time: N/A

Observation End Time: N/A


Just the status of the Musk Observatory was checked.

· Inside the Manual box was a battery (picture 1).

· Inside Quick Guides box, the Quick Guide and a hand control were found with an advice that said “Spare hand control. Please do not use unless instructed by the astronomy team”. Don’t worry, there is not an intention to use it (picture 2).

· The black box “Sirius Observatories” was turn on. After cheeked the full status of the Musk Observatory, I turned it off (picture 3).

· The picture of the astronomy box is attached (picture 4).

· The astronomy laptop was found in a case on the shelf in the lower hab (picture 5).

· In general, the Musk Observatory looks in good condition. I can’t wait to see the sun from Mars.

Objects Viewed: N/A

Problems Encountered: N/A

Journalist Report – January 15th

[Sol 03]

[The Force Awakens]

07:20 in the morning on the first floor of the habitat. Apparently someone woke up late but it didn´t stop us from doing the first exercise routine in days since we get Mars. Stretches, flexibility and some strength to start Sol 3. Then a breakfast made of orange and dried egg that looked like a yellow paste. The cooker promises to give it some more flavor tomorrow.

The day of the first Extra Vehicular Activity (EVA) for all Crew 187 arrived. Armed with our suits, helmets, oxygen tanks, gloves, boots and communication system the seven earthlings traveled to the nearby mountains and valleys with their vehicles to discover red lands and rocky horizons that make anyone sigh. Are we really on Mars? I wish everyone could see what we do.

When we walk outside it is impossible not to see in the sky those lines caused by the meteorites that cross the sky of the red planet that remind us of the stelae of the planes that fly in our home. Still nothing can be compared with the views that gave us the small Spark drone that traveled with us.

Then more than one hour outside for each expedition ended with a few minutes in the pressurization chamber and a pain in the back and shoulders of our XO that guided both EVAs. Objectives met. Success in the first EVA of Crew 187.

Hungry and tired we beg for food to our chefs who prepared rice with cheese and meat that reminded our Mexican crew members of a certain typical dish prepared with tortillas. Everyone asked for double ration but it was impossible because this will be our dinner.

After lunch we received a visit from the Habitat Director to conclude to understand and know the map of the area surrounding the habitat. There are many roads, many areas and many possibilities. We already have some expeditions in mind that will surely give us large amounts of information. Tomorrow four crew members will go to EVA to collect 15 kilograms of samples of analogue ground for their experiments.

With the night over us and about to conclude the activities we plan to have our first movies night with popcorn. It should be noted that get bored is not an option in the habitat with board games, space chess and more than 50 films for all tastes.

Those fans of space films will understand our excitement when we realize that a small Porg has nested in the EVA Officer’s bedroom. From now on, not only Earth life live in the habitat but also from planet Crait.

A sea of stars has arrived and now it covers us in this third night sleeping in our new home for another 13 nights surrounded by Martian red colour everywhere you see. LATAM II will continue to inform.

Tania Robles, MDRS Crew 187 Journalist

Picture of the day:

518800_425700 15Jan2018 Red Planet.jpg

Operations Report – January 15th

Crew 186 Operations Report 15012018

SOL: 03

Name of person filing report: L. Diaz

Non-nominal systems: None

Notes on non-nominal systems: Generator system limping along with a now-nominal routine.

Generator (hours run): 14h 54min

Generator turned off, charging battery at 9:30

Generator turned on at 8:26pm

Solar— SOC

@ 8:00am: 41%

@ 4:37pm: 93%

Diesel: 50%

Propane: 26%

Ethanol Free Gasoline (5 Gallon containers for ATV): 4.75 Gallons

Water (trailer): 125 Gallons

Water (static): 380 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Water Meter: 1296629 Gallons

Toilet tank emptied: No

ATVs Used: 350.1, 350.2, 350.3

Oil Added? No

ATV Fuel Used: – Gallons

# Hours the ATVs were used today: 00 hours 35 min

Notes on ATVs: ATVs were nominal.

Deimos rover used: No

Hours: –

Beginning charge: –

Ending charge:

Currently charging: No

Sojourner rover used: Assigned to director only.

Hours: Director discretional hours

Beginning charge:

Ending charge:

Currently charging: Maybe

Spirit rover used: No

Hours: –

Beginning charge: -%

Ending charge:

Currently charging: No

Opportunity rover used: No

Hours: –

Beginning charge:-

Ending charge:

Currently charging: No

Curiosity rover used: Yes

Hours: 35

Beginning charge: 100%

Ending charge:

Currently charging: Yes

HabCar used and why, where? No

General notes and comments: None

Summary of internet: All nominal

Summary of suits and radios: All nominal

Summary of Hab operations: All nominal

Summary of GreenHab operations: All nominal

Summary of ScienceDome operations: All nominal

Summary of RAM operations: Not Operational

Summary of health and safety issues: Add Rearview mirrors to the Rovers, is dangerous driving them without that.

Questions, concerns and requests to Mission Support: None

Greenhab Report – January 15th

GreenHab Report

Hernán David Mateus Jiménez


Environmental control:

Ambient with door opening

Shade cloth on

Working Hour: 6:15
Inside temp at working hour: 16° C
Outside temp during working hours: 2° C
Inside temperature high: 25° C
Inside temperature low: 15° C
Inside humidity: 73 %RH

Inside humidity high: 67 %RH
Inside humidity low: 23 %RH

Hours of supplemental light:

For the crops 05:00 to 11:59 PM

Changes to crops: The cucumber plants are a little wilted today

Daily water usage for crops: 8 gallons

Time(s) of watering for crops: 45 min

Research observations: The research has not started

Changes to research plants: The research has not started

Aquaponics: Not Functional

Narrative: Today I have had a great day, my first EVA on Mars, where I could test the necessary tools to collect Martian soil in my next EVA, after it, I did not realize that the temperature decrease, so I forgot to close the door at the afternoon, fortunately Atila did it. Also I had to repair some pieces of my project that were broken during the trip to Mars, so today I could not finish the assembly of the lysimeter.

Because of the fact that the source of water close to the greenhab is not working, I am spending a lot of time filling the watering can in the kitchen of the hab.

I hope to finish the assembly of the lysimeter and start the quinoa crop over it with the soil that we will get tomorrow at the EVA.

Sol Summary – January 15th

Crew 187 Sol 3 Summary Report 15JAN2018

Sol 3

Summary Title: Spaceman

Author’s name: Cynthia Fuertes Panizo

Mission Status: All nominal

Sol Activity Summary:

Mission Support,

Greetings from Mars!!! One more day in the red planet, it’s
breathtaking how every Martian has been integrated. At 07:20 we were together at the lower deck to do some relaxing exercises in the lead of by David, who also cleaned the Hab with Luis. At 08:00, we ate breakfast together while we did the briefing.

The day looked promising because today we will be our first EVAs. We wanted to assets the performance of the crew with the suits
operations, so the routine of each EVA was to walk over even terrain, then drive using ATVs and Curiosity; and finally we made summit in Overlook Reach. The first group of the EVA was conformed by Atila, Oscar Danton and I; their time of donning started at 11:00 until 11:30, in that moment the real adventure began and lasted until 13:00. Later, in the habitat they gave advices to the following group of EVA. The second group was conformed by Atila, Tania, Luis and David. Their expedition lasted until 15:30. Both EVA was led by Atila. After each EVA, the crew started to exchange their feelings about to be in a real expedition, that feeling of been a Martian is unique!

By the way, between the 2 EVAs, at 13:20, Tania and I, went to the Musk Observatory just to check the status.

As lunch, Oscar cooked rice with cheese and meal, it was a delicious lunch! Later, Shannon visited us to explain about the map on Mars. She suggested us to visit wonderful places like Copernicus highway, Lith canyon, Tank wash, Candor Chasma, and so on… We plan to visit those places, thanks for your advice!

Today we are going to have our first movie night in Mars, we are so excited for that!

Look Ahead Plan:

There are two main things that the crew are waiting for tomorrow, do an EVA to collect some examples for Atila and David’s experiments. In addition, the Musk Solar Observatory will be used… Can you imagine it? Look the sun from Mars!!!

Anomalies in work:

The commander room’s door was locked from inside. David and Atila worked together to open it again! Now we know that a locksmith is a profession in Mars.

Weather: Temperature: 2°C, Fair weather, Winds speed: NNE 6 MPH, Humidity 60%, Barometer 30.31in

Crew Physical Status: Healthy

Reports to be filed: Sol Summary, GeenHab Report, Operations Report, EVA Request, EVA Report, Astronomy Report, Journalist Report and HSO Report.

Support Requested: None.

EVA Report – January 15th

Purpose of EVA: Space Suit activity recognition and Assessment of in-suit performance


EVA 1: Atila, Danton, Oscar, Cynthia

EVA 2: Atila, Luis, David, Tania

Narrative: The crew conducted the first EVA set of activities today, and it was a huge success, whit all the objectives completed. The main goal of the EVA was to get all the crewmembers a first approach to external operations, and to test the main tools we would be using from now on, on each of the projects. The EVA was designed to also test the performance of the crewmembers with the gear on, to plan better the next EVAs, and assignments. Both crews performed the same routine, with small variations. The routine began by the donning of the suits, proceeding to enter the airlock, 5 minute depressurization, and going out in front of the hab. From there, the group took a straight route until the south of Marble Ritual, try the quadcopter operation, and the shovels for sampling. Upon return to the hab area, the crew left some of the equipment back and mounted the vehicles, 3 ATVs and the Curiosity Rover. It is important to clear that the requested Rover was Deimos, but the designated driver in EVA 1, Danton, would not fit inside of it, requiring the use of the one without safety cage. In order to keep the same protocols, EVA 2 used the same Rover. The next route was traversed by vehicle and took the crew to Pooh’s Corner, where a second brief walk was undertaken, towards a small mound on the East of Cow Dung Road. Afterwards the crew returned to the vicinity of the hab to perform the last test, which consisted of making summit of Overlook Ridge. The crew then returned to the airlock, sustained repressurization, and upon entry to the hab, proceeded to doffing protocols.

Each crewmember was interviewed upon finalization of the doffing on qualitative, and quantitative factors, as well as general comments on the performance of themselves on the suit. Further comment on this will be posted on the final report.

Oscar Ojeda

EVA Officer – Crew 187

Copyright © The Mars Society. All rights reserved. | Main Site